OpenStudy (anonymous):

Somone got counting formulas for me that i can prove? eg: tan(pi + X)= (1+tanx )/ (1-tanX)

OpenStudy (anonymous):

lhs=tan( pi+x)=tanx rhs=(1+tanx)/(1-tanx)=( 1+sinx/cosx)/(1- sinx/cosx)=cosx+sinx/cosx-sinx